161 Preliminary Report

Site 977

(Proposed site Alb-4A)

After the 99 nmi transit to Site 977 (~8 hr; ~12 nmi/hr), we conducted a seismic survey over the site. We then positioned over Site 977 and deployed a beacon at 2315 hr, 8 June, and a backup beacon at 0200 hr, 9 June. The elevation of the DES above sea level was 11.54 m for Hole 977A.

Hole 977A

We used the same APC/XCB BHA as had been run throughout the leg with a rerun security S87F bit. Since the site was in close proximity to a communication cable, the VIT was deployed as a precautionary measure while spudding Hole 977A. Hole 977A was spudded at 1100 hr, 9 June. Core 161-977A-1H was taken with the bit at 1990.0 mbrf and recovered 3.97 m; therefore, the seafloor was defined to be at 1984.0 mbsl. APC Cores 161-977A-1H to -17H were taken to 156.0 mbsf (2151.5 mbrf) and recovered 164.72 m (106%). Cores 161-977A-3H to -17H were oriented using the Tensor tool. ADARA temperature measurements were taken during Cores 161-977A-3H, -6H, -9H, and -12H. Some core expansion was observed beginning with Core 161-977A-12H. The maximum C1 detected in the upper 156 mbsf was 35,695 ppm, with a poorly defined second high of 31,564 at about 310 mbsf. No trends in the gas analyses were significant enough for us to consider abandoning the hole.

XCB Cores 161-977A-18X to -63X were taken from 156.0 to 598.5 mbsf (2151.5 to 2594.5 mbrf, respectively), coring 442.5 m and recovering 380.77 m (86.05%). From 156.0 to 531.0 mbsf, XCB recovery was over 100% in stiff, hard clay. Starting at Core 161-977A-57X (531.0 mbsf), we encountered a conglomeratic interval (cobble, gravel, and sand), and recovery from 531.0 to 598.5 mbsf fell to less than 4%. Prior to allowing the whole BHA to penetrate into this interval, we made a short wiper trip back up to 482.58 mbsf. Immediately after cutting the next core (Core 161-977A-63X), the pipe became stuck before the connection could be made and the core barrel withdrawn. We worked the pipe for over 2 hr until it finally became free. This zone acted as if it was slightly overpressured. When the mud pumps were turned off, the pipe was able to be worked free. It appeared that the added flow from the bit was causing the hole to pack off.

At this point, we decided to log instead of trying to fight the hole conditions, since the bit had already accumulated over 86 hr and we had the weaker nonmagnetic drill collar in the BHA. Therefore, we made a wiper trip to 49.5 mbsf (2045 mbrf) and positioned the bit at this depth for logging. Fill was tagged at 578.5 mbsf (2574 mbrf), but we did not attempt to ream it out due to the BHA conditions, described above, and the amount of time that might be lost if the drill pipe became stuck again. We ran the quad-combo and then the FMS. We did not run the GLT, since we had some difficulty reentering the drill pipe with the FMS tool. Both logs were run to 564.5 mbsf (2560 m) and will help provide information about the extremely low recovery interval below 531 mbsf.

After logging, we deployed an FFF. The VIT was used to observe the BHA being withdrawn from the seafloor to ensure that the FFF was not dragged out of the hole. The bit cleared the FFF at 0700 hr, 14 June. The RCB BHA was assembled and run to the seafloor. The BHA consisted of the same components that were used at earlier sites, except that the jars were not included since it was damaged during Hole 976E operations. It took about 15 min to locate and reenter the funnel. The FFF was reentered at 1800 hr, 14 June. Several attempts at raising and lowering the BHA were needed until the BHA found the original hole and could be lowered without bowing the pipe.

The pipe was lowered into the hole and encountered fill at 533.21 mbsf (2523.21 mbrf). We then picked up the top drive and easily washed and reamed to 595.18 mbsf (2590.68 mbrf) in less than 2 hr. The last 3 m of fill, however, could not be penetrated without the drill pipe becoming stuck. Every time the bit was gently placed on top of the fill, the torque increased to 600-700 amps, rotation stopped, and pressure increased. We suspect that the hole was packing off, since the only way to free the pipe was by completely turning off the mud pumps. We spent over 8 hr attempting to ream out this last 3 m, without success. During the last attempt, the weight indicator began fluctuating as if the bit were bouncing on the bottom. Since we were unable to advance, we decided to terminate the hole.

Prior to abandoning the hole, we requested to offset from this hole. We thought that moving out of the subsurface channel axis might allow us to penetrate past this interval to the 1200 mbsf target depth. However, these offset locations were denied for safety reasons. Instead, we received approval for drilling close to proposed site Alb-4 (slightly modified to be able to avoid drilling close to a submarine telecommunications cable). The bit cleared the seafloor at 1030 hr, and cleared the rotary at 1430 hr, 15 June. The two beacons were recovered while the pipe was being retrieved. After the BHA was recovered, we observed that one cone had broken off the new bit and that the welded collar around the location of the dogs on the MBR was heavily damaged. The rig floor was secured, and we began the transit to proposed site Alb-4B (Site 978) at 1430 hr, 15 June.

To Operations Site 978

161 Table of Contents

Publications Home

(ODP Home)